OpenStudy (gerald):

solve f(g)=sqtx^4-x^2

OpenStudy (anonymous):

What are you trying to solve for?

OpenStudy (anonymous):

If you mean\[f(g)=\sqrt{x^4}-x^2\]then that is 0, since\[\sqrt{x^4}=x^2\]

OpenStudy (anonymous):

Or do you mean,\[f(g)=(\sqrt{x})^4-x^2\]because that's 0 too since,\[(\sqrt{x})^4=(x^{\frac{1}{2}})^4=x^{4/2}=x^2\]